Near the end of a marathon race, the first two runners are separated by a distance of 75.0 m. The front runner has a velocityShow All work step by step. By what distance does the winning runner finish ahead of the next runner? distance:
Answer #1

75m2 9-4.75 e 215m Up 325 mis Finish line line misec VA-VB=4.75-3.25=1.5m/sec (2). Time taken by A to complete the race - Tim

C)

Time difference between Finishing of both runners = 3seconds ..

Relative velocity =1.5m/second

Distance when Winning runner ahead =1.5*3 =4.5m

75m2 9-4.75 e 215m Up 325 mis Finish line line misec VA-VB=4.75-3.25=1.5m/sec (2). Time taken by A to complete the race - Time 75+ 215 = 63.15 sec 4.75 Time sceken by B to complelee therace - Time = 215 = 66.15 sec 325 Second gunner will win the race. since (TB>TA)

At 6315 Sond, second semes, complete the pace, at 63 15 seconds, Font runner dun a distance 63.157325 = 205: 23m. Second runner run 3oom complete the sare while front runner needs 215m to complete But unable to cover, & 'cover only 205. 2bm 1 Intial condition OB (Frest sunnes) D 3oom OA att. (second runnen) Final condition 5 215m •-205:23m ( Second runna forest runner at time of time winner is abead of front Tu - 215-205.23 I 9.76ml